OpenStudy (anonymous):

Find the slope of a line perpendicular to 2x + y = -8.

OpenStudy (anonymous):

if two lines are perpndicular then product of thier slopes=-1 so let m is slope of given line and n is for desired (to be fouund) then m*n=-1 here m=-2, so -2n=-1 =>n=1/2

OpenStudy (haseeb96):

bring this equation into this form y=mx+b y=-2x-8 here the slope is m1=-2 so in the perpendicular liness condition is m1m2=-1 -2m2=-1 m2=-1/-2 m2=1/2 m=1/2 is the slope of perpendicular line